# Influence of initial states on memory effects: A study of early-time superradiance

### v1

@@ III.1 Theoretical model
The initial condition of the model is assumed to be $\rho_{S\!E}(t_{0})=\rho_{S}(t_{0})\otimes|0\rangle\langle 0|$ where $\rho_{S}(t_{0})$ is arbitrary and $|0\rangle$ is the vacuum state of the cavity. For simplicity, we use $t_{0}=0$ in the reminder of this paper without loss of generality. The density matrix of the composite system $\rho_{S\!E}$ is described by the master equation,
@@
$+\gamma(b\rho_{S\!E}b^{{\dagger}}-\frac{1}{2}b^{{\dagger}}b\rho_{S\!E}-\frac{1}{2}\rho_{S\!E}b^{{\dagger}}b)$
@@
In this remainder of this paper, we focus on the early-time dynamics where $gt\ll 1$ due to the following reasons. First, non-Markovian characters are non-negligible on such a short time scale. Second, it stresses the influence of a particular initial state since the state hardly changes in this time interval. Furthermore, it is helpful for understanding the creation mechanism of the superradiance. In the early-time limit $gt\rightarrow 0$, one might ignore the influence of cavity dissipation $\gamma$, providing that $\gamma$ is not infinitely large. In this case, $\rho_{S\!E}$ evolves unitarily via
$\rho_{S\!E}(t)=U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)$ (15)
@@ III.2 Early-time solution
$=$ $\sum_{k}\langle k|_{E}U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)|k\rangle_{E}$
@@
$e^{\alpha A}Be^{-\alpha A}=B+\alpha[A,B]+\frac{\alpha^{2}}{2!}[A,[A,B]]+...,$ (17)
after the replacement $A=H_{I}$, $B=\rho_{S}(0)\otimes\rho_{E}$ and $\alpha=-it$, we have
@@
where the terms with higher orders of $gt$ have been omitted due to the early-time limit ($gt\rightarrow 0$). Using the number state basis of the cavity $|k\rangle=|0\rangle,|1\rangle,...$ and the assumption $\rho_{E}=|0\rangle\langle 0|$, it is straightforward to derive the solution of the system evolution in the early-time limit $gt\rightarrow 0$,
@@
Here $\sigma^{-}=\sum_{n}\sigma_{n}^{-}$ is the collective lowering operator and $\mathcal{L}_{K}(\rho)=K\rho K^{+}-\frac{1}{2}K^{\dagger}K\rho-\frac{1}{2}\rho K^{\dagger}K$ is the Lindblad superoperator. $b^{{\dagger}}|0\rangle=|1\rangle$ and $\langle 0|b=\langle 1|$ have been used to obtain Eq.(19). Notice that the partial trace over the second term of Eq.(18) is zero, thus the change of $\rho_{S}(0)$ is quadratic in $t$ in the early-time limit.
@@ III.3 Memory effects
Eq.(26) demonstrates that in a short time interval $[0,t]$, $N_{M}[\rho(0)]$ grows quadratically with $t$. To focus on the influence of initial states (rather than the time $t$ or the atom-cavity coupling $g$), it is convenient to discuss the normalized value of memory effects
@@
It represents the strength of memory effects in the early-time limit as a function of only $\rho_{S}(0)$. Notice that $\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}$ is not bounded as $N_{M}[\rho_{S}(0)]$. In principle, there is $0\leqslant\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}<\infty$.
@@ III.4 Cavity photon number
In a non-Markovian proceess, the relation $\rho_{SE}(t)\approx\rho_{S}(t)\otimes\rho_{E}$ does not hold in general. Particularly, for the superradiance problem, the change of photo number in the environment (the cavity) might be an important source of the memory effects so that $\rho_{SE}(t)\neq\rho_{S}(t)\otimes\rho_{E}$. Thus it is desirable to know the cavity photon number (denoted by $N_{P}$ in this paper) in the early-time limit in order to understand the physics of the memory effects. The total excitation number of our model represented by $N_{ex}=b^{\dagger}b+\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}$ is conserved since $[N_{ex},H]=0$. Besides, there are no photon in the cavity initially. Therefore, the cavity photon number is equal to the loss of excitations of the atoms that represents the emission intensity of superradiance (in the early-time limit).
@@
It is observed that in the early-time limit, the cavity number increases quadratically with time. As mentioned in the last subsection, it is convenient to discuss the normalized cavity photon number
@@
such that it does not depend on the evolution time $t$ and atom-cavity coupling $g$. Eq.(30) represents the emission intensity in the early-time limit as a function of only the initial state. When each atom radiates independently, e.g., each atom radiates in its own cavity, the normalized cavity photon number for the $n$th atom is reduced to
$\frac{N_{P}^{(n)}[\rho_{S_{n}}(0)]}{(gt)^{2}}=\text{Tr}[\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S_{n}}(0)]$ (31)
where $\rho_{S_{n}}(0)$ is the reduced density matrix of the $n$th atom. Then the degree of superradiant might be measured by the ratio [7, 13]
$S[\rho_{S}(0)]$ $=$ $\frac{N_{P}[\rho_{S}(0)]/(gt)^{2}}{\sum_{n}N_{P}^{(n)}[\rho_{S}(0)]/(gt)^{2}}$ (32)
@@
for nonzero denominator. If $S[\rho_{S}(0)]$ is greater (less) than 1, the state $\rho_{S}(0)$ is superradiant (subradiant).
III.5 Early-time regime
Mathematically, Eq.(19) holds true as $gt\rightarrow 0$. One may wonder the validity of the early-time solution Eq.(19) in a longer time interval and how the cavity dissipation deteriorates the validity. In this subsection, we discuss this problem by comparing the systems dynamics by Eq.(19) with that by numerically solving the full dynamics in $\mathcal{H}_{S}\otimes\mathcal{H}_{E}$ with Eq.(14) and tracing out the environment. A dynamics could be represented by the dynamical maps $T(t,0)$ that turns all the possible initial state at 00 to a final state at $t$. Furthermore, the difference of two dynamical maps can be evaluated through their Choi-Jamiółkowski matrices [36, 37]. Here we evaluate the error of Eq.(19) at instant $t$ by the trace distance of $\rho_{T(t,0)}^{quad}$ and $\rho_{T(t,0)}^{exact}$, i.e.,
$Error=\frac{1}{2}\|\rho_{T(t,0)}^{quad}-\rho_{T(t,0)}^{exact}\|$ (33)
where $\rho_{T(t,0)}^{quad}$ and $\rho_{T(t,0)}^{exact}$ are the Choi-Jamiółkowski matrices calculated by Eq.(19) and Eq.(14), respectively. The Choi-Jamiółkowski matrix is defined as $\rho_{T}(t,0)=T(t,0)\otimes\mathbb{I}(|\Psi\rangle\langle\Psi|)$ with $|\Psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{i=1}^{2^{N}}|i\rangle|i\rangle$ a maximally entangled state of the $N$-atom system and an ancillary system of the same dimension.
We numerically calculated the error of Eq.(19) for $\gamma/g=0,1,10$ and $N=2,6$. The results are illustrated in Fig.2 in logarithmic scale. It is seen that the order of magnitude of error decreases linearly with that of $t$. Although the cavity dissipation and the increasing of atom number can increase the error, there exists a time interval where Eq.(19) is a good approximation as long as $\gamma$ and $N$ are finite. For example, in a time interval $[0,0.01/g]$, the error is less than $10^{-5}$ even when $\gamma=10g$ for both $N=2$ and $N=6$, implying that Eq.(19) is a good approximation in this scenario. The above discussion implies that the early-time limit $gt\rightarrow 0$ could be relaxed to an early-time regime, i.e., a time interval $[0,\tau]$ where the decay is quadratic. Here $\tau$ is linked to the Zeno time [38, 39, 40, 41].
@@ IV Influence of initial states
In this section, the results Eq.(27), Eq.(30) and Eq.(32) are applied to several types of $N$-atom initial states. Our first aim is to study the influence of initial states on the memory effects. Another aim is to reveal the role of memory effect in creating superradiance. Conversely, it also helps to understand the physical source of memory effects in our model.
@@ IV.1 Dicke states
$D_{z}|JM\rangle$ $=$ $M|JM\rangle\qquad(M=-J,...,J)$ (35)
where the operators are defined by $D^{+}=\sum_{n}\sigma_{n}^{+}$, $D^{-}=\sum_{n}\sigma_{n}^{-}$ and $D_{z}=\frac{1}{2}\sum_{n}(|e\rangle_{n}\langle e|_{n}-|g\rangle_{n}\langle g|_{n})$. A Dicke state is symmetrical and invariant by atom permutation with $N_{e}=J+M$ excited atoms and $N_{g}=J-M$ ground-state atoms. It can be constructed by the following formula [8]:
$|JM\rangle=\sqrt{\frac{(J+M)!}{N!(J-M)!}}(\sum_{n}\sigma_{n}^{-})^{J-M}|e,e,...,e\rangle.$ (36)
@@
A well-known result of the free-space spontaneous emission rate for two-level atoms is $W_{N}=\Gamma\langle\sigma^{+}\sigma^{-}\rangle$ where $\Gamma$ is the natural linewidth of a single atom. For the Dicke states, there is
@@
[…]
In the non-Markovian early-time regime, we calculated the normalized value of memory effects and cavity photon number with Eq.(27) and Eq.(30) for the Dicke states of different $N$. The results are shown in Fig.3 where $N=1,2,...,15$ ($J=N/2$) and $M=-J,-J+1,...,J$. It is observed that when the number of atoms $N$ is fixed, the strongest memory effects as well as photon emission happens when $M$ is 0 or next to 0, which is in agreement with Eq.(37). Interestingly, we find that the value of memory effects for the Dicke states can be fully determined by the cavity photon number by
$N_{M}(|JM\rangle)=\frac{1}{2}N_{P}(|JM\rangle)$ (38)
in the early-time regime. The relation is proved in the following. Using the property $\sigma^{\pm}|JM\rangle=\sqrt{J(J+1)-M(M\pm 1)}|J(M\pm 1)\rangle$, it is straightforward to see that
@@
where $M^{\prime}=M-1$. Remind that the trace norm could be calculated by $\|A\|=\text{Tr}\sqrt{A^{\dagger}A}=\sum_{m}|\lambda_{m}|$ where $\lambda_{m}$ is the eigenvalue of operator $A$. Meanwhile, the eigenvalues of $|JM^{\prime}\rangle\langle JM^{\prime}|-|JM\rangle\langle JM|$ is $1$ and $-1$ whose absolute values sum up to 2. Therefore,
@@
Considering that the normalized cavity photon number for Dicke states is
@@
We now discuss the degree of superadiance for the Dicke states in the early-time regime. The normalized cavity photon number for one independent atom in its excited state is given by $N_{P}(|e\rangle)/(gt)^{2}=N_{P}(|\frac{1}{2},\frac{1}{2}\rangle)/(gt)^{2}=1$ by Eq.(31). Using the results of Eq.(41), the degree of superradiance for the Dicke states could be calculated by
$S(|JM\rangle)$ $=$ $\frac{N_{P}(|JM\rangle)/(gt)^{2}}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)/(gt)^{2}}$ (42)
@@
for $J+M>0$, where the denominator represents the contribution of the emission from the $J+M$ independently excited atoms in $|JM\rangle$. Note that the physical meaning of the denominator in Eq.(42) is different from that in Eq.(32), but their values are the same.
In Fig.3(a) or (b), the left row represents $|J,-J\rangle$ (no atom excited), thus $N_{M}$ and $N_{P}$ are both zero which is a trivial case. Another group of Dicke states without superradiance are $|J,J\rangle$ (fully excited) corresponding to the right row in Fig.3(a) or (b). Using Eq.(42), it is easy to check that $S(|J,J\rangle)=1$. Except for the two rows, other Dicke states in Fig.3 are superradiant which could be examined by $S(|JM\rangle)>1$ with Eq.(42). For example, for the Dicke states $|J,-J+1\rangle$ (one atom excited) represented by the second row from the left, there is $S(|J,-J+1\rangle)=N$. The single-photon superradiance is called “the greatest radiation anomaly” by Dicke. Meanwhile, the normalized value of memory effects for $|J,-J+1\rangle$ is $\frac{N}{2}$, which is also proportional to $N$. For the states in the second row from the right, there is $S(|J,J-1\rangle)=2$ for $N>1$. The degree of superradiance is a constant showing a weaker degree of superradiance. For Dicke states, although the normalized value of memory effects is not proportional to the degree of superradiance in general, it is observed that the transition from non-supperradiant states to superradiant states (e.g., $|J,J\rangle\rightarrow|J,J-1\rangle$) is always accompanied by the increase of memory effects. Thus the memory effects are important for the creation of superradiance.

### v2

@@ III.1 Theoretical model
The initial condition of the model is assumed to be $\rho_{S\!E}(t_{0})=\rho_{S}(t_{0})\otimes|0\rangle\langle 0|$ where $\rho_{S}(t_{0})$ is arbitrary and $|0\rangle$ is the vacuum state of the cavity. For simplicity, we use $t_{0}=0$ in the remainder of this paper without loss of generality. The density matrix of the composite system $\rho_{S\!E}$ is described by the master equation,
@@
$+\gamma(b\rho_{S\!E}b^{{\dagger}}-\frac{1}{2}b^{{\dagger}}b\rho_{S\!E}-\frac{1}{2}\rho_{S\!E}b^{{\dagger}}b),$
@@
The superradiance phenomenon could be understood as the enhancement of collective spontaneous radiation caused by a common environment (compared with independent environments). To understand the role of memory effects in the superradiance process, we are also interested in the dynamics where the atoms radiate independently, e.g., each atom radiates in its own cavity. This dynamics could be described by the Hamiltonians $H_{I}=\sum_{n=1}^{N}g(\sigma_{n}^{+}b_{n}+\sigma_{n}^{-}b_{n}^{{\dagger}})$, $H_{E}=\sum_{n=1}^{N}\omega_{B}b_{n}^{{\dagger}}b_{n}$ and the environment initial state $\rho_{E}=\Pi_{n=1}^{N}|0\rangle_{n}\langle 0|_{n}$. Under the resonance condition, the master equation for the independently radiating atoms and the cavities in the interaction picture is given by
$\dot{\rho}_{S\!E}$ $=$ $-i[\sum_{n=1}^{N}g(\sigma_{n}^{+}b_{n}+\sigma_{n}^{-}b_{n}^{{\dagger}}),\rho_{S\!E}]$ (15)
$+\gamma\sum_{n=1}^{N}(b_{n}\rho_{S\!E}b_{n}^{{\dagger}}\!-\!\frac{1}{2}b_{n}^{{\dagger}}b_{n}\rho_{S\!E}\!-\!\frac{1}{2}\rho_{S\!E}b_{n}^{{\dagger}}b_{n}).$
III.2 Early-time solution
In this section, we focus on the early-time dynamics where $gt\ll 1$ due to the following reasons: First, non-Markovian characters are non-negligible on such a short timescale. Second, it stresses the influence of a particular initial state since the state hardly changes in this time interval. Furthermore, it is helpful for understanding the creation mechanism of the superradiance. In the early-time limit $gt\rightarrow 0$, one might ignore the influence of cavity dissipation $\gamma$, providing that $\gamma$ is not infinitely large. In this case, $\rho_{S\!E}$ evolves unitarily via
$\rho_{S\!E}(t)=U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t),$ (16)
@@ III.2 Early-time solution
$=$ $\sum_{k}\langle k|_{E}U(t)\rho_{S}(0)\otimes\rho_{E}U^{{\dagger}}(t)|k\rangle_{E},$
@@
$e^{\alpha A}Be^{-\alpha A}=B+\alpha[A,B]+\frac{\alpha^{2}}{2!}[A,[A,B]]+\cdots,$ (18)
after the replacement $A=H_{I}$, $B=\rho_{S}(0)\otimes\rho_{E}$, and $\alpha=-it$, we have
@@
where the terms with higher orders of $gt$ have been omitted due to the early-time limit ($gt\rightarrow 0$). Using the number state basis of the cavity $|k\rangle=|0\rangle,|1\rangle,\cdots$ and the assumption $\rho_{E}=|0\rangle\langle 0|$, it is straightforward to derive the solution of the system evolution in the early-time limit $gt\rightarrow 0$,
@@
Here $\sigma^{-}=\sum_{n}\sigma_{n}^{-}$ is the collective lowering operator and $\mathcal{L}_{K}(\rho)=K\rho K^{+}-\frac{1}{2}K^{\dagger}K\rho-\frac{1}{2}\rho K^{\dagger}K$ is the Lindblad superoperator. $b^{{\dagger}}|0\rangle=|1\rangle$ and $\langle 0|b=\langle 1|$ have been used to obtain Eq. (20). Notice that the partial trace over the second term of Eq. (19) is zero, thus the change of $\rho_{S}(0)$ is quadratic in $t$ in the early-time limit. Using similar procedures done above, we obtain the early-time solution for independently radiating atoms:
$\rho_{S}(t)=\rho_{S}(0)+(gt)^{2}\sum_{n=1}^{N}\mathcal{L}_{\sigma_{n}^{-}}[\rho_{S}(0)].$ (21)
@@ III.3 Memory effects
Equation (28) demonstrates that in a short time interval $[0,t]$, $N_{M}[\rho(0)]$ grows quadratically with $t$. To focus on the influence of initial states (rather than the time $t$ or the atom-cavity coupling $g$), it is convenient to discuss the normalized value of memory effects
@@
It represents the strength of memory effects in the early-time limit as a function of only $\rho_{S}(0)$. Notice that $\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}$ is not bounded as $N_{M}[\rho_{S}(0)]$. In principle, there is $0\leqslant\frac{N_{M}[\rho_{S}(0)]}{(gt)^{2}}<\infty$. Similarly, the normalized memory effects for independent cavities (denoted by $N_{M}^{ind}$) can be derived from Eq. (21). The result can be expressed simply by replacing $\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]$ in Eq. (29) by $\sum_{n=1}^{N}\mathcal{L}_{\sigma_{n}^{-}}[\rho_{S}(0)]$, i.e.,
$\frac{N_{M}^{\textrm{ind}}[\rho_{S}(0)]}{(gt)^{2}}=\frac{1}{4}\|\sum_{n=1}^{N}\mathcal{L}_{\sigma_{n}^{-}}[\rho_{S}(0)]\|.$ (30)
@@ III.4 Cavity photon number
In a non-Markovian proceess, the relation $\rho_{SE}(t)\approx\rho_{S}(t)\otimes\rho_{E}$ does not hold in general. Particularly, for the superradiance problem, the change of photon number in the environment (the cavity) might be an important source of the memory effects so that $\rho_{SE}(t)\neq\rho_{S}(t)\otimes\rho_{E}$. Thus it is desirable to know the cavity photon number (denoted by $N_{P}$ in this paper) in the early-time limit in order to understand the physics of the memory effects. The total excitation number of our model represented by $\hat{N}_{ex}=b^{\dagger}b+\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}$ is conserved since $[\hat{N}_{ex},H]=0$. Besides, there are no photon in the cavity initially. Therefore, the cavity photon number is equal to the loss of excitations of the atoms that represents the emission intensity of superradiance (in the early-time limit).
@@
It is observed that in the early-time limit, the cavity number increases quadratically with time. As mentioned in the last section, it is convenient to discuss the normalized cavity photon number
@@
such that it does not depend on the evolution time $t$ and atom-cavity coupling $g$. Equation (33) represents the emission intensity in the early-time limit as a function of only the initial state. When each atom radiates independently, e.g., each atom radiates in its own cavity, the normalized cavity photon number for the $n$th atom is reduced to
$\frac{N_{P}^{(n)}[\rho_{S_{n}}(0)]}{(gt)^{2}}=\text{Tr}[\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S_{n}}(0)],$ (34)
where $\rho_{S_{n}}(0)$ is the reduced density matrix of the $n$th atom. Then the degree of superradiant in the early-time regime might be measured by the ratio [9, 15]
$S[\rho_{S}(0)]$ $=$ $\frac{N_{P}[\rho_{S}(0)]}{\sum_{n}N_{P}^{(n)}[\rho_{S}(0)]}$ (35)
@@
for nonzero denominator. If $S[\rho_{S}(0)]$ is greater (less) than one, the state $\rho_{S}(0)$ is superradiant (subradiant).
III.5 Other manifestations of memory effects
As discussed in Sec. II, the difference between $\text{Tr}[P\rho(t_{2})]$ and $\text{Tr}[P\rho^{\prime}(t_{2})]$ can be used as an evidence or a manifestation of the memory effects. For the superradiance problem, the atom excitation number operator $P=\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}$ serves as a nature choice. Following similar treatments in Eqs. (26)-(28), it is seen that the maximum of the atom excitation difference $\Delta N_{ex}^{\textrm{atom}}[\rho_{S}(0)]=|\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho_{S}(t_{2})]-\text{Tr}[\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rho^{\prime}_{S}(t_{2})]|$ in the time interval $[0,t]$ also happens at $\tau_{10}=\tau_{21}=\frac{t}{2}$ and
$\max_{\tau_{10},\tau_{21}}\Delta N_{ex}^{\textrm{atom}}[\rho_{S}(0)]$ (36)
$=$ $|2g^{2}\text{Tr}\{\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}(\frac{t}{2})^{2}\mathcal{L}_{\sigma^{-}}[\rho_{S}(0)]\}|.$
Using the results given by Eq. (31) and (32), the manifestation can be further simplified by
$\max_{\tau_{10},\tau_{21}}\Delta N_{ex}^{\textrm{atom}}[\rho_{S}(0)]$ $=$ $\frac{(gt)^{2}}{2}\text{Tr}[\sigma^{+}\sigma^{-}\rho_{S}(0)]$ (37)
$=$ $\frac{1}{2}N_{P}[\rho_{S}(0)].$
It is seen that the maximal difference of the atom excitation numbers in $[0,t]$ is exactly half the cavity photon number at $t$ in the early-time limit. Therefore, the nonzero cavity photon number itself is also a manifestation of memory effects.
III.6 Early-time regime
Mathematically, Eq. (20) holds true as $gt\rightarrow 0$. One may wonder the validity of the early-time solution Eq. (20) in a longer time interval and how the cavity dissipation deteriorates the validity. In this section, we discuss this problem by comparing the systems dynamics by Eq. (20) with that by numerically solving the full dynamics in $\mathcal{H}_{S}\otimes\mathcal{H}_{E}$ with Eq. (14) and tracing out the environment. A dynamics could be represented by the dynamical maps $T(t,0)$ that turns all the possible initial states at 00 to their final states at $t$. Furthermore, the difference of two dynamical maps can be evaluated through their Choi-Jamiółkowski matrices [38, 39]. Here we evaluate the error of Eq. (20) at instant $t$ by the trace distance of $\rho_{T(t,0)}^{\textrm{quad}}$ and $\rho_{T(t,0)}^{\textrm{exact}}$, i.e.,
$Error=\frac{1}{2}\|\rho_{T(t,0)}^{\textrm{quad}}-\rho_{T(t,0)}^{\textrm{exact}}\|$ (38)
where $\rho_{T(t,0)}^{\textrm{quad}}$ and $\rho_{T(t,0)}^{\textrm{exact}}$ are the Choi-Jamiółkowski matrices calculated by Eq. (20) and (14), respectively. The Choi-Jamiółkowski matrix is defined as $\rho_{T}(t,0)=T(t,0)\otimes\mathbb{I}(|\Psi\rangle\langle\Psi|)$ with $|\Psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{i=1}^{2^{N}}|i\rangle|i\rangle$ a maximally entangled state of the $N$-atom system and an ancillary system of the same dimension.
@@ III.6 Early-time regime
Figure 3: The error of Eq. (20) in terms of the atom excitation number difference for (a) 2 and (b) 6 fully excited atoms. Other parameters are the same with those in Fig. 2.
We numerically calculated the error of Eq. (20) for $\gamma/g=0,1,10$ and $N=2,6$. The results are illustrated in Fig. 2 in logarithmic scale. It is seen that the order of magnitude of error decreases linearly with that of $t$. Although the cavity dissipation and the increasing of atom number can increase the error, there exists a time interval where Eq. (20) is a good approximation as long as $\gamma$ and $N$ are finite. For example, in a time interval $[0,0.01/g]$, the error is less than $10^{-5}$ even when $\gamma=10g$ for both $N=2$ and $N=6$, implying that Eq. (20) is a good approximation in this scenario. The above discussion implies that the early-time limit $gt\rightarrow 0$ could be relaxed to an early-time regime, i.e., a time interval $[0,\tau]$ where the decay is quadratic. Here $\tau$ is linked to the Zeno time [40, 41, 42, 43]. We also illustrate the difference of the atom excitation numbers from Eqs. (20) and (14), i.e., $Error_{ex}=|\langle\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rangle_{quad}-\langle\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\rangle_{exact}|$, in Fig. 3 as another manifestation of the error. The initial states in Figs. 3(a) and 3(b) are 2 and 6 fully excited atoms, respectively. The result is similar to that in Fig. 2 but easier to be verified experimentally.
@@ IV Influence of initial states
In this section, the results Eqs. (29), (30), (33) and (35) are applied to several types of $N$-atom initial states. Our first aim is to study the influence of initial states on the memory effects. Another aim is to reveal the role of memory effect in creating superradiance. Conversely, it also helps to understand the physical source of memory effects in our model.
@@ IV.1 Dicke states
$D_{z}|JM\rangle$ $=$ $M|JM\rangle\qquad(M=-J,\cdots,J)$ (40)
where the operators are defined by $D^{+}=\sum_{n}\sigma_{n}^{+}$, $D^{-}=\sum_{n}\sigma_{n}^{-}$, and $D_{z}=\frac{1}{2}\sum_{n}(|e\rangle_{n}\langle e|_{n}-|g\rangle_{n}\langle g|_{n})$. A Dicke state is symmetrical and invariant by atom permutation with $N_{e}=J+M$ excited atoms and $N_{g}=J-M$ ground-state atoms. It can be constructed by the following formula [10]:
$|JM\rangle=\sqrt{\frac{(J+M)!}{N!(J-M)!}}(\sum_{n}\sigma_{n}^{-})^{J-M}|e,e,\cdots,e\rangle.$ (41)
@@
A well-known result of the free-space spontaneous emission rate for two-level atoms is $W_{N}=\Gamma\langle\sigma^{+}\sigma^{-}\rangle$, where $\Gamma$ is the natural linewidth of a single atom. For the Dicke states, there is
@@
Figure 4: (a) Normalized memory effects $N_{M}$ and (b) normalized cavity photon number $N_{P}$ in the early-time regime for different Dicke states $|JM\rangle$. The number of atoms $N$ varies from 1 to 15, correspondingly, $J=N/2=1/2,3/2,\cdots,15/2$. The results demonstrate that if the atoms are initially in $|JM\rangle$, the normalized value of memory effects is half the normalized cavity photon number in the early-time regime. The strongest memory effects as well as photon emission happens when $M$ is 0 or next to 0.
In the non-Markovian early-time regime, we calculated the normalized value of memory effects and cavity photon number with Eqs. (29) and (33) for the Dicke states of different $N$. The results are shown in Fig. 4 where $N=1,2,\cdots,15$ ($J=N/2$) and $M=-J,-J+1,\cdots,J$. It is observed that when the number of atoms $N$ is fixed, the strongest memory effects as well as photon emission happens when $M$ is 0 or next to 0, which is in agreement with Eq. (42). Interestingly, we find that the value of memory effects for the Dicke states can be fully determined by the cavity photon number in the early-time regime:
$N_{M}(|JM\rangle)$ $=$ $\frac{1}{2}N_{P}(|JM\rangle).$ (43)
Considering the result in Eq. (37), there is also
$N_{M}(|JM\rangle)=\max_{\tau_{10},\tau_{21}}\Delta N_{ex}^{\textrm{atom}}[|JM\rangle].$ (44)
Thus the value of memory effects can be directly measured through the atom excitation number difference or the cavity photon number. The relation Eq. (43) is proved in the following. Using the property $\sigma^{\pm}|JM\rangle=\sqrt{J(J+1)-M(M\pm 1)}|J(M\pm 1)\rangle$, it is straightforward to see that
@@
where $M^{\prime}=M-1$. Remind that the trace norm could be calculated by $\|A\|=\text{Tr}\sqrt{A^{\dagger}A}=\sum_{m}|\lambda_{m}|$, where $\lambda_{m}$ is the eigenvalue of operator $A$. Meanwhile, the eigenvalues of $|JM^{\prime}\rangle\langle JM^{\prime}|-|JM\rangle\langle JM|$ is $1$ and $-1$ whose absolute values sum up to 2. Therefore,
@@
Considering that the normalized cavity photon number for the Dicke states is
@@
Figure 5: (a) The degree of superradiance $S$ and (b) the degree of memory-effect-enhancement $N_{M}/N_{M}^{\textrm{ind}}$ (from independent radiation to collective radiation) in the early-time regime for different Dicke states. The states are the same with those in Fig. 4 except that $S$ and $N_{M}/N_{M}^{\textrm{ind}}$ for $M=-J$ are undefined and not shown. For a given $J$, the highest degree of superradiance happens when $M=-J+1$ (single excitation).
We now discuss the degree of superadiance for the Dicke states in the early-time regime. The normalized cavity photon number for one independent atom in its excited state is given by $N_{P}(|e\rangle)/(gt)^{2}=N_{P}(|\frac{1}{2},\frac{1}{2}\rangle)/(gt)^{2}=1$ by Eq. (34). Using Eq. (47), the degree of superradiance for the Dicke states could be calculated by
$S(|JM\rangle)$ $=$ $\frac{N_{P}(|JM\rangle)}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)}$ (48)
@@
for $J+M>0$ as shown in Fig. 5(a). Here the denominator represents the contribution of the emission from the $J+M$ independently excited atoms in $|JM\rangle$. Note that the physical meaning of the denominator in Eq. (48) is different from that in Eq. (35), but their values are the same. In Fig. 5(a), $S$ for $|J,-J\rangle$ (no atom excited) are undefined and not shown. Another group of Dicke states without superradiance are $|J,J\rangle$ (fully excited) corresponding to the right row in Fig. 5(a). Except for the two groups, other Dicke states are superradiant which could be examined by $S(|JM\rangle)>1$ with Eq. (48). For example, for $|J,-J+1\rangle$ (one atom excited) represented by the left row in Fig. 5(a), there is $S(|J,-J+1\rangle)=N$. The single-photon superradiance is called “the greatest radiation anomaly” by Dicke. For the states in the right row in Fig. 5(a), there is $S(|J,J-1\rangle)=2$ for $N>1$, showing a weaker degree of superradiance.
The degree of superradiance represents the magnification of the radiation intensity caused by a common environment (compared with $N$ independent environments). One interesting question is whether the memory effects are simultaneously enhanced by a common environment. To uncover the role of memory effects in the superradiance process, we next examine the degree of memory-effect-enhancement, i.e., $N_{M}(|JM\rangle)/N_{M}^{\textrm{ind}}(|JM\rangle)$, where $N_{M}^{\textrm{ind}}(|JM\rangle)$ is the value of memory effects with $N$ independent cavities calculated by Eq. (30). We find by mathematical induction that the value of memory effects for independently radiating atoms in a Dicke state is $N_{M}^{\textrm{ind}}(|JM\rangle)=\frac{1}{2}(gt)^{2}(J+M)$. Therefore, the degree of memory-effect-enhancement in the early-time regime is
$\frac{N_{M}(|JM\rangle)}{N_{M}^{\textrm{ind}}(|JM\rangle)}=J-M+1=S(|JM\rangle)$ (49)
as shown in Fig. 5(b). The enhancement of radiation intensity (caused by a common environment) is equal to the enhancement of memory effects for the Dicke states, implying that the memory effects of a common environment contribute to the establishment of superradiance.